OpenStudy (anonymous):

What is HX(1/sqrt(2) |0> + 1/sqrt(2) |+>)?

OpenStudy (anonymous):

What is \[HX\left(\frac{1}{\sqrt(2)}|0\rangle+\frac{1}{\sqrt(2)}|+\rangle\right)\] ?

OpenStudy (anonymous):

Can someone show me step by step?

OpenStudy (turingtest):

remember that\[|0\rangle=\binom10\]and\[|1\rangle=\binom01\]

OpenStudy (turingtest):

X is the bit-flip gate\[X=\begin{bmatrix} 0&1\\1&0\end{bmatrix}\]so what is\[X\otimes|0\rangle\]and\[X\otimes|1\rangle\]?

OpenStudy (turingtest):

oh I see the other guy is plus. yeah, you should rewrite it first I guess

OpenStudy (turingtest):

you will have to renormalize it

OpenStudy (turingtest):

no I'm sorry, you should separate them

OpenStudy (anonymous):

yes I think I stuck at the normalizing part.....

OpenStudy (anonymous):

I think I draw faster.